# Uniform pathwise stability of additive singular SDEs driven by fractional Brownian motion

### v1

@@ 4 Existence and uniqueness of invariant measures
  • $\llbracket X-B\rrbracket_{\mathcal{C}^{\alpha}_{I}L_{m,q}}$ is replaced by $\llbracket\Psi(t_{0},\cdot,x,\eta)-\widetilde{B}^{t_{0}}-\eta\rrbracket_{\mathcal{C}^{\alpha}_{I}L_{m,q}}=\llbracket\Psi(t_{0},\cdot,x,\eta)-\widetilde{B}^{t_{0}}\rrbracket_{\mathcal{C}^{\alpha}_{I}L_{m,q}}$.
  • $X-B]_{\mathcal{C}^{\alpha}_{I}L_{m}}$ is replaced by $\Psi(t_{0},\cdot,x,\eta)-\widetilde{B}^{t_{0}}-\eta]_{\mathcal{C}^{\alpha}_{I}L_{m}}$.
Those results will be summarised in Theorem 4.2, they are to be compared to Theorem 3.1, Theorem 3.2, Proposition 3.10 and Proposition 3.12. Their proof is omitted and the dependence of constants on $\eta$ is clarified. In particular, recall that the uniform-in-time bound on the moments goes through a comparison with the Ornstein-Uhlenbeck process, now it will go through a comparison with the process
$dU(t_{0},t,x,\eta)=-U(t_{0},t,x,\eta)dt+d\eta_{t}+d\widetilde{B}^{t_{0}}_{t},\quad U(t_{0},t_{0},x,\eta)=x.$
In particular, one can show (as in the proof of Proposition 3.8) that the $m$-th moment of $\Psi(t_{0},t,x,\eta)-U(t_{0},t,x,\eta)$ is bounded as follows:
$\|\Psi(t_{0},t,x,\eta)-U(t_{0},t,x,\eta)\|_{L_{m}}\leq C_{1}(1+C_{0}(\kappa,T-t_{0})\sup_{r\in[t_{0},T]}\|U(t_{0},t,x,\eta)\|_{L_{m}})$
where $C_{0}$ is given by (3.23) and $C_{1}$ does not depend on time. Notice that $U(t_{0},\cdot,x,\eta)$ is a Gaussian process of bounded variance, and its mean $m(t_{0},t,x,\eta)$ is given by
$m(t_{0},t,x,\eta):=e^{-(t-t_{0})}x+\eta_{t}-e^{-(t-t_{0})}\eta_{t_{0}}-\int_{t_{0}}^{t}e^{-(t-u)}\eta_{u}du.$
Hence, there exists a constant $C=C(m,H,x)$ such that for any $t\geq 0$,
$\|U(t_{0},t,x,\eta)\|_{L_{m}}\leq C(m(t_{0},t,x,\eta)+1).$
Therefore, the bound on the moments of $\Psi$ in the following theorem will depend on $\eta$ through $m(t_{0},t,x,\eta)$.
@@
Let $m\in[2,\infty)$ and $\gamma\in(1-1/(2H),1)$. Let $b$ satisfy Assumption 2.4. Let $F$ satisfy (2.11) and further assume that $F$ satisfies Assumption 2.3. Let $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$, $t_{0}<T$, and $\eta:[t_{0},T]\rightarrow\mathbb{R}^{d}$ be a continuous function, then there exists a unique solution $(\Psi(t_{0},t,x,\eta))_{t\in[t_{0},T]}$ to (4.1) in the class
$\widetilde{\mathcal{V}}^{\eta}:=\{Y:\,[Y-\widetilde{B}^{t_{0}}-\eta]_{\mathcal{C}^{1/2}_{[t_{0},T]}L_{2}}<\infty\}.$ (4.6)
Moreover, the following holds
@@
    for any $m\in[2,\infty)$, there exist constants $C_{1}:=C_{1}(m,\gamma,H,x,\kappa,\Xi)$ and $C_{2}:=C_{2}(m,\gamma,H,\kappa,\Xi)$ such that for any $t_{0}\leq s<t$ with $t-s\leq 1$, we have
    $\|\Psi(t_{0},t,x,\eta)\|_{L_{m}}$ $\leq C_{1}(1+\sup_{r\in[t_{0},T]}m(t_{0},r,x,\eta))\Big(1+(T-t_{0})\wedge\frac{1-e^{-\kappa_{1}(T-t_{0})}}{\kappa_{1}}\Big),$
@@
    $[\Psi(t_{0},\cdot,x,\eta)-\widetilde{B}^{t_{0}}-\eta]_{\mathcal{C}^{1+H(\gamma\wedge 0)}_{[s,t]}L_{m}}$ $\leq C_{2}(1+\sup_{r\in[s,t]}\|\Psi(t_{0},r,x,\eta)\|_{L_{m}}).$
@@
    there exists a constant $C:=C(m,\gamma,H,L_{F},\Xi)$ such that for any $t_{0}\leq s<t$ with $t-s\leq 1$, we have
    $\llbracket\Psi(t_{0},\cdot,x,\eta)-\widetilde{B}^{t_{0}}\rrbracket_{\mathcal{C}^{1+H(\gamma\wedge 0)}_{[s,t]}L_{m,\infty}}\leq C.$ (4.7)
@@
    Let $g$ satisfy Assumption 2.4, and denote by $\Psi_{g}$ the unique solution to (4.1) with singular drift $g$ in the class $\widetilde{\mathcal{V}}^{\eta}$, then there exists a constant $C:=C(m,\gamma,H,L_{F},\Xi,T-t_{0})$ such that for all $t\in[t_{0},T]$
    $\|\Psi(t_{0},t,x,\eta)-\Psi_{g}(t_{0},t,x,\eta)\|_{L_{m}}\leq C\|b-g\|_{\mathcal{C}^{\gamma-1}}.$ (4.8)
  4. (iv)
    If $\kappa_{2}=0$, then there exists a constant $\mathbf{M}=\mathbf{M}(m,\gamma,H,L_{F},\Xi)$ such that for any $t\in[t_{0},T]$, $\Psi$ satisfies
    $\|\Psi(t_{0},t,x,\eta)-\Psi(t_{0},t,y,\eta)\|_{L_{m}}\leq|x-y|\exp\{-\kappa_{1}+\mathbf{M}\|b\|_{\mathcal{C}^{\gamma}}(1+\|b\|_{\mathcal{C}^{\gamma}})^{\frac{4}{1+H(\gamma-1)}}(t-t_{0})\}.$ (4.9)
The following theorem shows that the solutions of (1.1) and (4.1) are the same when $\eta=\bar{B}^{t_{0}}$. A solution of (1.1) will be understood with respect to the filtered probability space and Wiener process fixed before.
@@
Let $\gamma\in(1-1/(2H),1)$ and let $b$ satisfy Assumption 2.4. Let $F$ satisfy Assumption 2.3 and further assume that $F$ satisfies (2.11) with $\kappa_{2}=0$. For any $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$, $t_{0}<T$, and any continuous function $\eta:[t_{0},T]\rightarrow\mathbb{R}^{d}$, denote by $(\Psi(t_{0},t,x,\eta))_{t\in[t_{0},T]}$ and $(X^{x}_{t})_{t\in[t_{0},T]}$ the respective unique solutions to (4.1) and (1.1) in the respective classes $\widetilde{\mathcal{V}}^{\eta}$ and $\mathcal{V}$ starting from initial condition $x$. Then for any $t\in[t_{0},T]$, $(\eta,\omega)\mapsto\Psi(t_{0},t,x,\eta)(\omega)$ has a jointly measurable modification, which we still call $\Psi$, and almost surely for $\omega\in\Omega$, we have
$X^{x}_{t}(\omega)=\Psi(t_{0},t,x,\bar{B}^{t_{0}}(\omega))(\omega).$ (4.10)
The proof of Theorem 4.3 is established in Section 4.1. The final theorem deals with the existence and uniqueness of invariant measures when $\kappa_{1}>0$.
Theorem $4.$
Let $m\in[2,\infty)$. Let $\gamma\in(1-1/(2H),1)$, and let $b$ satisfy Assumption 2.4. Let $F$ satisfy Assumption 2.3 and further assume that $F$ satisfies (2.11). Then the following statements hold:
@@
    If $\kappa_{1}>0$ in (2.11), then there exists an invariant measure associated to (1.1).
@@
    Recall the constant $\beta(m,\|b\|_{\mathcal{C}^{\gamma}})$ defined in (3.5). If $\kappa_{1}>0$, $\kappa_{2}=0$ and the $\mathcal{C}^{\gamma}$-norm of $b$ is sufficiently small so that $\beta(2,\|b\|_{\mathcal{C}^{\gamma}})<0$, then there exists a unique invariant measure associated to the SDE (1.1).
@@
4.1 Link with the SDE driven by the whole fBm: Proof of Theorem 4.3
Before moving on to proving existence and uniqueness of an invariant measure, the goal of this section is to prove Theorem 4.3 which describes the link between the SDE (4.1) we are using to define an invariant measure and the original SDE (1.1). In order to compare the two SDEs, in this section, the solution to (1.1) will be understood with respect to the filtered probability space and Wiener process fixed at the beginning of Section 4. Before comparing the two solutions, we need an intermediary result regarding the stability of (4.1) with respect to the history of the fBm.
@@
Let $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$, $t_{0}<T$ and $\eta^{1},\eta^{2}:[t_{0},T]\rightarrow\mathbb{R}^{d}$ be two continuous functions. Let $F$ satisfy Assumption 2.3 and further assume that $F$ satisfies (2.11) with $\kappa_{2}=0$. Let $\gamma\in(1-1/(2H),1)$ and let $b$ satisfy Assumption 2.4. Let $(\Psi(t_{0},t,x,\eta^{1}))_{t\in[t_{0},T]},(\Psi(t_{0},t,x,\eta^{2})_{t\in[t_{0},T]}$ denote the two unique solutions, in the respective classes $\widetilde{\mathcal{V}}^{\eta^{1}},\widetilde{\mathcal{V}}^{\eta^{2}}$, to the SDE (4.1) with initial condition $x$, respective paths $\eta^{1},\eta^{2}$, and singular drift $b$.
Let $m\in[2,\infty)$. There exists a constant $C:=C(m,\gamma,H,L_{F},\Xi,T-t_{0})$ such that for all $t\in[t_{0},T]$, the following holds
$\|\Psi(t_{0},t,x,\eta^{1})-\Psi(t_{0},t,x,\eta^{2})\|_{L_{m}}\leq C\sup_{t\in[t_{0},T]}|\eta^{1}_{t}-\eta^{1}_{t_{0}}-(\eta^{2}_{t}-\eta^{2}_{t_{0}})|.$ (4.11)

### v2

@@ 4 Markovian structure and invariant measures
  • $\llbracket X-B\rrbracket_{\mathcal{C}^{\alpha}_{I}L_{m,q}}$ is replaced by $\llbracket X^{(X_{0},V)}(t_{0},\cdot)-\widetilde{B}^{t_{0}}-V\rrbracket_{\mathcal{C}^{\alpha}_{I}L_{m,q}}=\llbracket X^{(X_{0},V)}(t_{0},\cdot)-\widetilde{B}^{t_{0}}\rrbracket_{\mathcal{C}^{\alpha}_{I}L_{m,q}}$.
  • $X-B]_{\mathcal{C}^{\alpha}_{I}L_{m}}$ is replaced by $X^{(X_{0},V)}(t_{0},\cdot)-\widetilde{B}^{t_{0}}-V]_{\mathcal{C}^{\alpha}_{I}L_{m}}$.
Moreover, since the innovation $(\widetilde{B}^{t_{0}}_{t})_{t\geq t_{0}}$ is independent of $\mathcal{F}_{t_{0}}$, the well-posedness analysis of Section 3 still holds for an $\mathcal{F}_{t_{0}}$ random variable $(X_{0},V)$. This holds for example if $X_{0}$ is $\mathcal{F}_{t_{0}}$-measurable and $V=\mathbf{A}^{t_{0}}(\bar{W})$ where $\bar{W}$ is a continuous random process such that $\bar{W}_{t}$ is $\mathcal{F}_{t_{0}}$-measurable for any $t\leq 0$. Well-posedness results for the SDE (4.2) will be summarised in Theorem 4.1, they are to be compared to Theorem 3.1, Theorem 3.2, Proposition 3.10 and Proposition 3.12. Their proof is omitted, and the dependence of constants on $V$ is clarified. In particular, recall that the uniform-in-time bound on the moments goes through a comparison with the Ornstein-Uhlenbeck process, now it will go through a comparison with the process
$dU^{(X_{0},V)}(t_{0},t)=-U^{(X_{0},V)}(t_{0},t)dt+d\eta_{t}+d\widetilde{B}^{t_{0}}_{t},\quad U^{(X_{0},V)}(t_{0},t_{0})=X_{0}.$
In particular, one can show (as in the proof of Proposition 3.8) that the $m$-th moment of $X^{(X_{0},V)}(t_{0},t)-U^{(X-0,V)}(t_{0},t)$ is bounded as follows:
$\|X^{(X_{0},V)}(t_{0},t)-U^{(X_{0},V)}(t_{0},t)\|_{L_{m}}\leq C(1+C_{0}(\kappa,T-t_{0})\sup_{r\in[t_{0},T]}\|U^{(X_{0},V)}(t_{0},r)\|_{L_{m}})$
where $C_{0}$ is given by (3.23). Notice that $U^{(X_{0},V)}(t_{0},\cdot)$ is a Gaussian process with bounded variance, and its mean $m^{(X_{0},V)}(t_{0},t)$ is given by
$m^{(X_{0},V)}(t_{0},t):=e^{-(t-t_{0})}\mathbb{E}(X_{0})+\mathbb{E}(V_{t})-e^{-(t-t_{0})}\mathbb{E}(V_{t_{0}})-\int_{t_{0}}^{t}e^{-(t-u)}\mathbb{E}(V_{u})du.$
Hence, there exists a constant $C=C(m,H,x)$ such that for any $t\geq t_{0}$,
$\|U^{(X_{0},V)}(t_{0},t)\|_{L_{m}}\leq C(|m^{(X_{0},V)}(t_{0},t)|+1).$
Therefore, the bound on the moments of $X^{(X_{0},V)}$ in the following theorem will depend on $V$ through $m^{(X_{0},V)}(t_{0},t)$.
@@
Let $m\in[2,\infty)$ and $\gamma\in(1-1/(2H),1)$. Let $b$ satisfy Assumption 2.2, $F$ satisfy Assumptions 2.3 and 2.4. Let $T\in\mathbb{R}$, $t_{0}<T$, $X_{0}:\Omega\rightarrow\mathbb{R}^{d}$ be an $\mathcal{F}_{t_{0}}$-measurable random variable, and $V:\Omega\rightarrow\mathcal{C}([t_{0},+\infty),\mathbb{R}^{d})$ be a random variable such that $V_{t}$ is $\mathcal{F}_{t_{0}}$-measurable for any $t\geq t_{0}$. Then there exists a unique solution $(X^{(X_{0},V)}(t_{0},t))_{t\in[t_{0},T]}$ to (4.2) in the class
$\mathcal{V}^{V}:=\{Y:\,[Y-\widetilde{B}^{t_{0}}-V]_{\mathcal{C}^{1/2}_{[t_{0},T]}L_{2}}<\infty\}.$ (4.8)
Moreover, the following statements hold.
@@
    There exist $C_{1}:=C_{1}(m,\gamma,H,\kappa,\Xi)$, $C_{2}:=C_{2}(m,\gamma,H,\kappa,\Xi)$ and $C_{3}:=C_{3}(m,\gamma,H,L_{F},\Xi)$ such that for any $t_{0}\leq s<t$ with $t-s\leq 1$, we have
    $\|X^{(X_{0},V)}(t_{0},t)\|_{L_{m}}$ $\leq C_{1}(1+\sup_{r\in[t_{0},T]}|m^{(X_{0},V)}(t_{0},r)|)\Big(1+(T-t_{0})\wedge\frac{1-e^{-\kappa_{1}(T-t_{0})}}{\kappa_{1}}\Big),$ (4.9)
    $[X^{(X_{0},V)}(t_{0},\cdot)-\widetilde{B}^{t_{0}}-V]_{\mathcal{C}^{1+H(\gamma\wedge 0)}_{[s,t]}L_{m}}$ $\leq C_{2}(1+\sup_{r\in[s,t]}\|X^{(X_{0},V)}(t_{0},r)\|_{L_{m}}),$
@@
    $\llbracket X^{(X_{0},V)}(t_{0},\cdot)-\widetilde{B}^{t_{0}}\rrbracket_{\mathcal{C}^{1+H(\gamma\wedge 0)}_{[s,t]}L_{m,\infty}}\leq C.$ (4.10)
@@
    Let $g$ satisfy Assumption 2.2, and denote by $X_{g}^{(X_{0},V)}$ the unique solution to (4.2) in the class $\mathcal{V}^{V}$, with initial input $(X_{0},V)$ and singular drift $g$, then there exists a constant $C:=C(m,\gamma,H,L_{F},\Xi,T-t_{0})$ such that for all $t\in[t_{0},T]$
    $\|X^{(X_{0},V)}(t_{0},t)-X_{g}^{(X_{0},V)}(t_{0},t)\|_{L_{m}}\leq C\|b-g\|_{\mathcal{C}^{\gamma-1}}.$ (4.11)
@@
    If $\kappa_{2}=0$, then there exists a universal constant $C$ and a constant $\mathbf{M}=\mathbf{M}(m,\gamma,H,L_{F},\Xi)$ such that for any $\mathcal{F}_{t_{0}}$-measurable random variable $Y_{0}$ and any $t\in[t_{0},T]$, we have
    $\begin{split}&\|X^{(X_{0},V)}(t_{0},t)-X^{(Y_{0},V)}(t_{0},t)\|_{L_{m}}\\ &\leq\|X_{0}-Y_{0}\|_{L_{m}}\exp\{(-\kappa_{1}+\mathbf{M}\|b\|_{\mathcal{C}^{\gamma}}(1+\|b\|_{\mathcal{C}^{\gamma}})^{\frac{4}{1+H(\gamma-1)}})(t-t_{0})\}.\end{split}$ (4.12)
The well-posedness assumptions of Theorem 4.1 will always be assumed in the following results. From now on, for any $t_{0}\in\mathbb{R}$, for any $\mathcal{F}_{t_{0}}$-measurable random variable $(X_{0},V):\Omega\rightarrow\mathbb{R}^{d}\times\mathcal{C}([t_{0},+\infty),\mathbb{R}^{d})$, $X^{(X_{0},V)}(t_{0},\cdot)$ will denote the unique solution to (4.2) in the class $\mathcal{V}^{V}$. The following theorem proves the existence of the map $(x,v)\mapsto X(t_{0},t,x,v)$, hinted at in (4.3), for any $t\geq t_{0}$ and states that $\Theta$ defined in (4.6) is a Markov process and justifies defining an invariant measure as a generalised initial condition that satisfies $P^{*}_{t}\mu=\mu$ for all $t\geq 0$. This is a classical result when $b$ is smooth. Here the challenge lies in proving the result for singular $b$. First, we recall the definition of evolution families and time non-homogeneous Markov processes.
Definition 4.2.
Let $t_{0}\in\mathbb{R}$ and $E$ be a Polish space, and denote by $\mathcal{B}_{b}(E)$ the space of bounded measurable functions $f:E\rightarrow\mathbb{R}$.
We say that $P:=(P_{s,t})_{t_{0}\leq s\leq t}$ is a evolution family on $E$ if for any $t_{0}\leq s\leq u\leq t$ $P_{s,t}$ maps $\mathcal{B}_{b}(E)$ to $\mathcal{B}_{b}(E)$, $P_{s,s}$ is the identity operator and $P_{s,u}P_{u,t}=P_{s,t}$. We say that $P:=(P_{s,t})_{t_{0}\leq s\leq t}$ is a Feller evolution family if it is an evolution family and if for any $t_{0}\leq s\leq t$, $P_{s,t}$ maps $\mathcal{C}(E)$ to $\mathcal{C}(E)$.
Given an evolution family $P$, we say that $(Y_{t})_{t\geq t_{0}}$ is a Markov process with evolution family $P$ if for any $t_{0}\leq s\leq t$ and any $f\in\mathcal{B}_{b}(E)$, we have
$\mathbb{E}^{s}f(Y_{t})=P_{s,t}f(Y_{s})\text{ a.s}.$
@@
Let $\gamma\in(1-1/(2H),1)$. Let $b$ satisfy Assumption 2.2 and $F$ satisfy Assumptions 2.3 and 2.4. Then the following statements hold.
@@
    For any $t_{0}\in\mathbb{R}$ and any $t\geq t_{0}$, there exists a jointly-measurable map $(x,v)\mapsto X(t_{0},t,x,v)$ such that for all $(x,v)\in\mathbb{R}^{d}\times\mathcal{C}([t_{0},\infty),\mathbb{R}^{d})$,
    $\mathbb{P}(X^{(x,v)}(t_{0},t)=X(t_{0},t,x,v))=1.$
@@
    For any $t_{0}\in\mathbb{R}$ and any $(x,w)\in\mathbb{R}^{d}\times\mathcal{H}_{H}$, define the evolution $\Theta:=(\Theta(t_{0},t,x,w))_{t\geq t_{0}}$ as in (4.6) and the family of operators $P:=(P_{s,t})_{t_{0}\leq s\leq t}$ as in (4.7). Then, $P$ is a Feller evolution family on $\mathbb{R}^{d}\times\mathcal{H}_{H}$, and $\Theta$ is a Markov process with evolution family $P$.
  3. (iii)
    For any $t_{0}\in\mathbb{R}$, any $t_{0}\leq s\leq t$ and any generalised initial condition $\mu_{0}$, we have that $P^{*}_{s,t}\mu_{0}=P^{*}_{0,t-s}\mu_{0}$, and $(P_{r}^{*}\mu_{0})_{r\geq 0}:=(P^{*}_{0,r}\mu_{0})_{r\geq 0}$ forms a semigroup, where for any $r\geq 0$, $P^{*}_{0,r}$ denotes the adjoint of the operator $P_{0,r}$.
The proof of Theorem 4.3 is established in Section 4.1. We now provide a rigorous definition of an invariant measure.
Definition 4.4.
Under the assumptions of Theorem 4.3, we say that $\mu$ is an invariant measure associated to (4.2) if $\mu$ is a generalised initial condition and satisfies $P^{*}_{t}\mu=\mu$ for any $t\geq t_{0}$.
The final theorem deals with the existence and uniqueness of invariant measures when $\kappa_{1}>0$.
Theorem 4.5.
Let $\gamma\in(1-1/(2H),1)$ and let $b$ satisfy Assumption 2.2. Let $F$ satisfy Assumptions 2.3 and 2.4 . Then the following statements hold.
  1. (i)
    If $\kappa_{1}>0$ in (2.12), then there exists an invariant measure associated to (4.2) that has finite moments.
  2. (ii)
    Let $m\in[2,\infty)$ and recall the constant $\beta(m,\|b\|_{\mathcal{C}^{\gamma}})$ defined in (3.5). If $\kappa_{1}>0$, $\kappa_{2}=0$ and the $\mathcal{C}^{\gamma}$-norm of $b$ is sufficiently small so that $\beta(2,\|b\|_{\mathcal{C}^{\gamma}})<0$, then there exists a unique invariant measure $\mu$ associated to the SDE (4.2). Moreover, there exists a constant $C:=C(m,\gamma,H,\kappa,\Xi)$ such that for any $t_{0}\in\mathbb{R}$, any $t\geq t_{0}$ and any $\mathcal{F}_{t_{0}}$-measurable random variable $(X_{0},\bar{W})$ whose law is a generalised initial condition, we have
    $\mathcal{W}_{m}(\mathcal{L}(\Theta^{(X_{0},\bar{W})}(t_{0},t),\mu)\leq C\exp\{\beta(m,\|b\|_{\mathcal{C}^{\gamma}})(t-t_{0})\}.$ (4.13)
@@
4.1 Markov property: Proof of Theorem 4.3
Before moving on to proving existence and uniqueness of the invariant measure, the goal of this section is to prove Theorem 4.3 which motivates the definition of the invariant measure. We will always assume that $F$ satisfies Assumptions 2.3 and 2.4, $\gamma\in(1-1/(2H),1)$ and $b$ satisfies Assumption 2.2. Therefore, existence and uniqueness of solutions to (1.1) (in the class (3.4)) and (4.2) (in the class (4.8)) will follow from Theorem (3.2) and Theorem (4.1). We start by proving the joint-measurability of the map $(x,v,\omega)\mapsto X^{(x,v)}(t_{0},t)(\omega)$ in Lemma 4.7. The proof is based on the following proposition, which gives the continuity of $X^{(x,v)}(t_{0},t)$ with respect to $v$.
@@
Let $x\in\mathbb{R}^{d}$, $v^{1},v^{2}\in\mathcal{C}([t_{0},+\infty),\mathbb{R}^{d})$, $T\in\mathbb{R}$ and $t_{0}<T$. For any $m\in[2,\infty)$, there exists a constant $C:=C(m,\gamma,H,L_{F},\Xi,T-t_{0})$ such that for all $t\in[t_{0},T]$, the following holds
$\|X^{(x,v^{1})}(t_{0},t)-X^{(x,v^{2})}(t_{0},t)\|_{L_{m}}\leq C\sup_{r\in[t_{0},T]}|v^{1}_{r}-v^{1}_{t_{0}}-(v^{2}_{r}-v^{2}_{t_{0}})|.$ (4.14)